Find the value(s) of d so that the points (6,d) and (4,7) lie on a line that is parallel to the line 7x-8y=-80
Answer by HyperBrain(694)

7x-8y=-80
---
We transform this line into its slope-intercept form.
7x-8y=-80
8y=7x+80
y=(7/8)x+10
---
Now, we know that the slope of this line is 7/8.
The line passing through the given two points must also have a slope of 7/8 for it to be parallel to the given line.
---
(6,d) and (4,7)
m=(d-7)/(6-4)=(d-7)/2
(d-7)/2=7/8
4(d-7)=7
4d-28=7
4d=35
d=35/4